7.8 Given the income elasticity of money demand is 0.6, and the interest elasticity of money demand is –0.3, if Y increases by 5% AND the interest rate falls by 15% of what it was (e.g., the interest rate decreases from 10.0% to 8.5%), then what would the central bank have to do to keep prices stable (i.e., to keep   0).
Answer #1

Y increases by 5%

Thus, increase in money demand due to increase in Y = 0.6*5% = 3%

Also, increase in money demand due to decrease in interest rate = 0.3*15% = 4.5%.

Thus, total increase in money demand = 7.5%

Therefore, to keep prices stable, the central bank has to increase money supply by 7.5% to match up the increase in money demand.

Thanks!